Jody invested $5000 less in an account paying 4% simple interest than she did in an account paying 3% simple interest. At the end of the first year, the total interest from both accounts was $675. Find the amount invested in each amount invested in each account.
Answer by checkley77(12569)

.04(X-5,000)+.03X=675
.04X-200+.03X=675
.07X=675+200
.07X=875
X=875/.07
X=12,500 AMOUNT INVESTED @ 3%
12,500-5,000=7,500 INVESTED @ 3%
PROOF:
.04*7,500+.03*12,500=675
300+375=675
675=675